OpenStudy (anonymous):

sin^(-1)⁡(sin⁡〖-135°〗 ) solve for theta.

OpenStudy (anonymous):

The answer is an angle measure. But I am confused how to get it. It's 135 or -135 I believe but I need to know how to simplify it.

OpenStudy (anonymous):

I'm getting -pi/4, not 135. Maybe I am misunderstanding the question.

OpenStudy (anonymous):

I think that is the problem that I am having. I believe that the answer is supposed to be an angle measure.

OpenStudy (anonymous):

the instructions say "evaluate the following. answers must be in exact values"

OpenStudy (anonymous):

-45 in degrees, again, not 135.

OpenStudy (anonymous):

I just want to see how you derived that. I know that it is a 45 degree reference angle and that the -45 degrees is the same as -135 but I need to show it. That is where I am confused.

OpenStudy (anonymous):

pi/4 = 45 degrees

OpenStudy (anonymous):

ahh, 1 min..

OpenStudy (anonymous):

\[\sin^{-1} (\cos(135\deg+\frac{ \pi }{ 2}))\]